Grade 12th passAlgebra

Simplify and solve the following linear equations.
3(5z – 7) – 2(9z – 11) = 4(8z – 13) – 17

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

3(5z–7)–2(9z–11)=4(8z–13)–17
⇒15z-21-18z+22=32z-52-17
⇒15z-18z-32z=-52-17+21-22
⇒-35z=-70
⇒z = -70/-35
⇒z=2

Thanks